# Intrinsic spin distributions in multinucleon transfer reactions

### v1

@@ 2 Theoretical framework
[…]
where $\Omega=(\alpha,\beta,\gamma)$ represents the three Euler angles. $D^{S}_{MK}(\Omega)$ is the Wigner function, where $M$ and $K$ are projections of the angular momentum $\vec{S}$ on the symmetry axis in the laboratory and intrinsic frames, respectively. $\hat{R}(\Omega)$ is the rotation operator:
@@
with the intrinsic spin operators ${\hat{S}}_{x,y,z}$ defined in the center-of-mass frame of the respective fragment. The probability distribution for the PLF or TLF with an intrinsic spin $S$ and projection $K$ is computed from the expression:
@@
Since the angular momentum operator is a one-body operator, the matrix $\langle\mathrm{\Psi}|e^{-i\alpha\hat{S}_{z}}e^{-i\beta\hat{S}_{y}}e^{-i\gamma\hat{S}_{z}}|\mathrm{\Psi}\rangle$ can be expressed through rotation of single-particle states $|\psi_{k}\rangle$
@@
For the stationary solution, we use a 3D lattice with dimensions $N_{x}\times N_{y}\times N_{z}=26\times 26\times 26$ and a mesh spacing of 0.8 fm, while dynamical calculations employ an asymmetric lattice of $60\times 26\times 60$ with the same spacing. The Dirac equation (Eq. (2)) is solved numerically using a predictor-corrector method with a time step of 0.2 fm/$c$. The initial configuration positions the projectile and target nuclei 24 fm apart on the lattice, assuming Rutherford trajectories prior to collision. The time evolution continues until either two well-separated fragments emerge with a center-of-mass distance exceeding 20 fm or 24 fm (impact-parameter dependent), or the system evolves into a composite nucleus (fusion). In angular momentum projection, for the integration over angles, we employ trapezoidal integration with uniform discretized points in the $\alpha$ and $\gamma$ directions, while using non-uniform Gaussian integration in the $\beta$ direction. The Euler angles are discretized with $30\times 30\times 30$ mesh points for the PLF and $40\times 40\times 40$ for the TLF, ensuring sufficient resolution for spin distribution analysis. To obtain the projection along the $K$-axis, which is defined as the line connecting the centers of mass of the PLF and TLF, we rotate the nuclei to let the $K$-axis be parallel to the $z$-axis before projecting on angular momentum.
@@ 3 Results and discussion
Taking the final fragments of the reactions at impact parameters $b=6$ fm ($\Lambda=109.9~{}\hbar$), $5$ fm ($\Lambda=91.6~{}\hbar$), and $4.65$ fm ($\Lambda=85.2~{}\hbar$) as examples, the intrinsic spin distributions of the PLF and the TLF are shown in Fig. 1. Since the symmetry with respect to the $x$-$z$ plane, for both fragments $P(S,K)=P(S,-K)$, only the distributions of $K\geq 0$ are displayed here. At $b\geq 6$ fm, 40Ca and 208Pb barely come into contact, and the average number of transferred nucleons is close to zero [58]. As shown in Fig. 1(a) and (d), the spins of both fragments are nearly zero.
@@
To assess the validity of entanglement Shannon entropy as a measure of correlations between two observables, we first examine the relationship between $S$ and $|K|$, noting the symmetry $P(S,K)=P(S,-K)$. Figure 2 displays the Shannon entropies derived from the distributions $P(S,|K|)$, $P(|K|)$, and $P(S)$ in the PLF, along with the mutual information. The marginal distributions of the $S$ and $|K|$ are obtained via $P(S)=\sum_{K}P(S,K)$ and $P(|K|)=\sum_{S}P(S,|K|)$. At large impact parameters (e.g., $b=10$ fm), the entanglement Shannon entropy is nearly zero, as evidenced by $P(S=0,|K|=0)=P(S=0)\times P(|K|=0)=1$, indicating negligible correlation. However, as the impact parameter decreases, the entanglement Shannon entropy rises to approximately $1.5$, reflecting enhanced correlations between $S$ and $|K|$. This trend is further supported by the fragment spin distributions at $b=5$ fm and $b=4.65$ fm, shown in Fig. 1(b) and (c). If $S$ and $|K|$ were independent, the joint distribution would factorize as $P(S,|K|)=P(S)P(|K|)$, subject to the geometric constraint $S\geq|K|$. However, the distributions in Figs. 1(b) and (c) deviate from this expectation, indicating a correlation between $S$ and $|K|$. In Fig. 2, the mutual information initially follows the trend of the Shannon entropies, increasing for smaller impact parameters $b$. However, for $b<6$ fm, $I(S,|K|)$ no longer mirrors the Shannon entropies and even exhibits a slight decline. A similar trend is observed for the corresponding quantities in the TLF. This suggests that as nucleon transfer between the nuclei intensifies, the correlation between $S$ and $|K|$ weakens. As shown in Fig. 1(b) and (c), while more orbital angular momentum is converted into fragment spins, their projections along the symmetry axis ($|K|$) remain confined to relatively small values. This implies that the spins are predominantly oriented perpendicular to the collision plane. Consequently, the mutual information $I(S,|K|)$ does not increase with the Shannon entropies at smaller impact parameters.
@@
[…]
where $k=K_{\text{PLF}}$ ($K_{\text{TLF}}$) is the spin projection along a certain direction for the PLF (TLF).
We analyze two principal directions, as illustrated schematically in the upper right panel of Fig. 3: (1) the $K$-axis defined as the line connecting the centers of mass of the two fragments and, (2) the $y$-axis, perpendicular to the reaction plane. The distributions of the intrinsic spin projections along the $K$- and $y$-axes are presented in Figs. 3 and 4, respectively.

### v3

@@ 2 Theoretical framework
$\hat{P}^{S}_{MK}=\frac{2S+1}{16\pi^{2}}\int d\Omega(D^{S}_{MK}(\Omega))^{*}\hat{R}(\Omega),$ (4)
where $\Omega=(\alpha,\beta,\gamma)$ represents the three Euler angles. $D^{S}_{MK}(\Omega)$ is the Wigner function, where $M$ and $K$ are projections of the intrinsic spin $\vec{S}$ on the symmetry axis in the laboratory and body-fixed frames, respectively. $\hat{R}(\Omega)$ is the rotation operator:
@@
with ${\hat{S}}_{y,z}$ representing the intrinsic spin operators along $y$- and $z$-axis. The probability distribution for the PLF or TLF with an intrinsic spin $S$ and projection $K$ is computed from the expression:
@@
Since the intrinsic spin operator is a one-body operator, the overlap $\langle\mathrm{\Psi}|e^{-i\alpha\hat{S}_{z}}e^{-i\beta\hat{S}_{y}}e^{-i\gamma\hat{S}_{z}}|\mathrm{\Psi}\rangle$ can be expressed through rotation of single-particle states $|\psi_{k}\rangle$
@@
For the stationary solution, we use a 3D lattice with dimensions $N_{x}\times N_{y}\times N_{z}=26\times 26\times 26$ and a mesh spacing of 0.8 fm, while dynamical calculations employ an asymmetric lattice of $60\times 26\times 60$ with the same spacing. The Dirac equation (Eq. (2)) is solved numerically using a predictor-corrector method with a time step of 0.2 fm/$c$. The initial configuration positions the projectile and target nuclei 24 fm apart on the lattice, assuming Rutherford trajectories prior to collision. The time evolution continues until either two well-separated fragments emerge with a center-of-mass distance exceeding 20 fm or 24 fm (impact-parameter dependent), or the system evolves into a composite nucleus (fusion). In angular momentum projection, for the integration over angles, we employ trapezoidal integration with uniform discretized points in the $\alpha$ and $\gamma$ directions, while using non-uniform Gauss-Legendre integration in the $\beta$ direction. The Euler angles are discretized with $30\times 30\times 30$ mesh points for the PLF and $40\times 40\times 40$ for the TLF, ensuring sufficient resolution for spin distribution analysis. To obtain the projection along the $K$-axis, which is defined as the line connecting the centers of mass of the PLF and TLF, we rotate the nuclei to let the $K$-axis be parallel to the $z$-axis before projecting on angular momentum.
@@ 3 Results and discussion
Taking the final fragments of the reactions at impact parameters $b=6$ fm ($\Lambda=109.9$$\hbar$), 55 fm ($\Lambda=91.6$$\hbar$), and $4.65$ fm ($\Lambda=85.2$$\hbar$) as examples, the intrinsic spin distributions of the PLF and the TLF are shown in Fig. 1. Due to the $x$-signature symmetry, the following relation holds: P(S,K)=P(S,K)P(S,-K)=P(S,K) and we display only the distributions for $K\geq 0$. At $b\geq 6$ fm, 40Ca and 208Pb barely come into contact, and the average number of transferred nucleons is close to zero [58]. As shown in Fig. 1(a) and (d), the spins of both fragments are nearly zero.
@@
To assess the validity of entanglement Shannon entropy as a measure of correlations between two observables, we first examine the relationship between $S$ and $|K|$, noting the symmetry P(S,K)=P(S,K)P(S,K)=P(S,-K). Figure 2 displays the Shannon entropies derived from the distributions $P(S,|K|)$, $P(|K|)$, and $P(S)$ in the PLF, along with the mutual information. The marginal distributions of the $S$ and $|K|$ are obtained via $P(S)=\sum_{K}P(S,K)$ and $P(|K|)=\sum_{S}P(S,|K|)$. At large impact parameters (e.g., $b=10$ fm), the entanglement Shannon entropy is nearly zero, as evidenced by $P(S=0,|K|=0)=P(S=0)\times P(|K|=0)=1$, indicating negligible correlation. However, as the impact parameter decreases, the entanglement Shannon entropy rises to approximately $1.5$, reflecting enhanced correlations between $S$ and $|K|$. This trend is further supported by the fragment spin distributions at $b=5$ fm and $b=4.65$ fm, shown in Fig. 1(b) and (c). If $S$ and $|K|$ were independent, the joint distribution would factorize as $P(S,|K|)=P(S)P(|K|)$, subject to the geometric constraint $S\geq|K|$. However, the distributions in Figs. 1(b) and (c) deviate from this expectation, indicating a correlation between $S$ and $|K|$. In Fig. 2, the mutual information initially follows the trend of the Shannon entropies, increasing for smaller impact parameters $b$. However, for $b<6$ fm, $I(S,|K|)$ no longer mirrors the Shannon entropies and even exhibits a slight decline. A similar trend is observed for the corresponding quantities in the TLF. This suggests that as nucleon transfer between the nuclei intensifies, the correlation between $S$ and $|K|$ weakens. As shown in Fig. 1(b) and (c), while more orbital angular momentum is converted into fragment spins, their projections along the symmetry axis ($|K|$) remain confined to relatively small values. This implies that the spins are predominantly oriented perpendicular to the collision plane. Consequently, the mutual information $I(S,|K|)$ does not increase with the Shannon entropies at smaller impact parameters.
@@
where the symbol $(\text{P})$ or $(\text{T})$ denotes the projection operator that only acts on the wave function of the PLF or TLF, respectively, defined as:
$\hat{P}_{k}=\frac{1}{4\pi}\int_{0}^{4\pi}d\theta e^{i\theta k}e^{-i\theta\hat{S}_{m}},$ (13)
where $\hat{S}_{m}$ is the intrinsic spin operator along a certain $m$-axis, $k=K_{\text{PLF}}$ ($K_{\text{TLF}}$) is the spin projection along the $m$-axis for the PLF (TLF).
We analyze two principal directions, as illustrated schematically in the upper right panel of Fig. 3: (1) the $K$-axis defined as the line connecting the centers of mass of the two fragments, and (2) the $y$-axis, perpendicular to the reaction plane. The distributions of the intrinsic spin projections along the $K$- and $y$-axes are presented in Figs. 3 and 4, respectively.